A cannonball is fired horizontally at 23.40 m/s from the edge of a 55 m building. Determine the: (a) height of the cannonball from the ground after 2 s. (b) time taken by the cannonball to reach the ground. (c) distance from the base of the building to the point where the its landed.


1
Expert's answer
2022-01-24T13:15:57-0500

Given:

"v=\rm 23.40\: m/s"

"H=\rm55\: m"

"g=\rm 9.8\:m/s^2"


(a) height of the cannonball from the ground after 2 s

"h=H-gt^2/2=55-9.8*2^2/2=35.4\:\rm m"

(b) time taken by the cannonball to reach the ground

"t=\sqrt{\frac{2H}{g}}=\sqrt{\frac{2*55}{9.8}}=3.35\: \rm s"

(c) distance from the base of the building to the point where the its landed

"L=vt=23.40*3.35=78.4\;\rm m"